OpenStudy (anonymous):

Jason considered two similar televisions at a local electronics store. The generic version was based off the brand name and was 3/4 the size of the brand name. If the generic television set is 12 inches by 24 inches, what are the dimensions of the brand name television? A)18 inches by 36 inches B)6 inches by 12 inches C)16 inches by 32 inches D)9 inches by 18 inches

OpenStudy (anonymous):

Almost identical to this: http://openstudy.com/study#/updates/4ff7e7abe4b01c7be8ca9a19 That explanation applies here.

OpenStudy (anonymous):

do 3/4 divided by 12 and 3/4 divided by 24 and you will get both of your answers

OpenStudy (anonymous):

iss this on flvs? lol who is your teacher? im doing the same test right now lol

OpenStudy (anonymous):

yeah it is ! idk her name lol

OpenStudy (anonymous):

No, you wouldn't do (3/4)/12. That would get you (3/48)

OpenStudy (anonymous):

Sidenote: FLVS students: Are you actually in FL, or one of the myriad of students across the world?

OpenStudy (anonymous):

actually in florida !

OpenStudy (anonymous):

i got A ?

OpenStudy (anonymous):

that gets you 9 wired.. 3 divided by 4 equals 0.75 that times 12 equals 9.

OpenStudy (anonymous):

the answer is d

OpenStudy (anonymous):

alright im redoing it

OpenStudy (anonymous):

@Nyole you said "3/4 divided by 12" @softballlover6655 Not A, as that would be if it was (2/3).

OpenStudy (anonymous):

oh sorry. guess im sleepy :o wired can you help me with the last question i posted

OpenStudy (anonymous):

Generic TV = 3/4ths of the Brand Name Generic = (3/4) Brand Multiply both sides by (4/3) to eliminate the fraction on the right: (4/3) Generic = Brand

OpenStudy (anonymous):

ah thanks !

OpenStudy (anonymous):

No problem. What did you get for the answer?

OpenStudy (anonymous):

12*(4/3) by 24*(4/3) =16 by 32

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

Yep!